# Adversarial Attacks and Defenses in Automated Control Systems: A Comprehensive Benchmark

### v2

@@ Abstract
Integrating machine learning into Automated Control Systems (ACS) enhances decision-making in industrial process management. One of the limitations to the widespread adoption of these technologies in industry is the vulnerability of neural networks to adversarial attacks. This study explores the threats in deploying deep learning models for fault diagnosis in ACS using the Tennessee Eastman Process dataset. By evaluating three neural networks with different architectures, we subject them to six types of adversarial attacks and explore five different defense methods. Our results highlight the strong vulnerability of models to adversarial samples and the varying effectiveness of defense strategies. We also propose a novel protection approach by combining multiple defense methods and demonstrate it’s efficacy. This research contributes several insights into securing machine learning within ACS, ensuring robust fault diagnosis in industrial processes.
@@ 1 Introduction
Automated Control Systems (ACS) operate with a variety of digital and analog signals received from sensors and control mechanisms. The conventional paradigm involves processing and analyzing this data using simplistic rules embedded in programmable logic controllers. However, the limitations of such systems become evident when dealing with complex signal combinations, impeding precise diagnosis of technological process states. To address this challenge, there is currently a shift towards the development of machine learning algorithms in ACS [1]. Modern neural networks can significantly improve industrial process management, but face a substantial limitation — their vulnerability to adversarial attacks.
The purpose of this study is to analyze the problem of adversarial attacks on deep learning models implemented in ACS. Moreover, our objective is to explore effective methodologies to mitigate these challenges. As an example of ACS, we use the Tennessee Eastman Process dataset [2] where the task is the Fault Detection and Diagnosis (FDD) in a chemical process.
Fault Detection and Diagnosis methods, also known as techniques for identifying and diagnosing faults, play a pivotal role in monitoring diverse industrial processes, ranging from chemical processes to electromechanical drive systems. The classification proposed by [3], categorizes these methods into three groups — those based on expert knowledge, mathematical models, and data-driven approaches. The latter one includes various approaches in machine learning, including neural networks.
In this work, we explore how adversarial attacks and defenses against them can affect neural networks with different architectures. We consider three various deep learning models — Multi-Layer Perceptron (MLP), model based on Gated Recurrent Units (GRU), and Temporal Convolutional Network (TCN). We subject these models to six different types of adversarial attacks and explore five defense methods. A novel protection strategy is proposed, which consists of combining several defense methods. We then analyze the success of these protective measures [^1] The source code to reproduce our results is available at https://github.com/AIRI-Institute/fdd-defense..
Our contributions are the following:
  • We find that different defenses are effective to varying degrees against different attacks. Universal defense methods greatly reduce models quality on original data.
  • We propose a new protection strategy as a combination of several defense methods and demonstrate its average effectiveness against various attacks.
@@ 2.1 Classification of attacks on machine learning models
Attacks on machine learning models are typically categorized into several types, which vary based on the capabilities of the attacker in relation to the target and its characteristics throughout the model’s lifecycle. These attack types include evasion attacks [5, 6, 7], poison attacks [8, 4, 9] and exploratory attacks [10, 11, 12]. Additionally, some of these attack categories have further subgroups. In this article, we specifically examine evasion attacks. Furthermore, attacks are segmented into three groups based on the level of information available about the model’s architecture and access to its internal parameters and/or requests: white-box, black-box, and gray-box attacks.
@@
In white-box attacks [13, 14, 15], the attacker possesses complete information, enabling them to execute any operations on their instance of the deployed model (e.g., obtaining gradients, accessing output data from any layer) to construct perturbed samples. In the presence of defense mechanisms, these mechanisms are also susceptible to attack by the adversary.
Black-box attacks [16, 17, 18, 19] assume that the attacker can only make a limited number of requests ($L$, where $1\leq L<\infty$) to the deployed model. The models provide the attacker with predictions such as label class or probability, semantic map segmentation, etc.
Gray box (or semi-white box) [20, 21, 22] attacks represent an intermediate state between white-box and black-box scenarios. They involve the imposition of certain restrictions that provide some information about the learning model/process, albeit incomplete.
@@ 2.2 Most common methods of attack and defense
In the literature, there are numerous articles that explore issues akin to those addressed in this study. However, these methods are either examined on different datasets or pertain to disparate domain areas, necessitating adaptation for our purposes. Notably, many initial attacks were devised for image analysis models, and not all methods have been fully tailored for the domain area under our investigation. Therefore, in this review, we also consider these aspects. Articles such as [23, 24, 25, 26, 27] delve into the impact of attacks on various image datasets like CIFAR-10, CIFAR-100, ResNet-20, and MNIST. These articles also discuss various protective measures. Key attack methods include L-BFGS, FGSM, PGD, C&W, and DeepFool. While identifying the most prevalent defense methods can be challenging, several key approaches emerge, notably Defense-GAN and Adversarial Training. Moreover, articles such as [28, 29, 30, 31, 32, 33, 34, 35] discuss attacks and defense methods on datasets like TEP and/or similar domain areas such as CARLA, Electra, SWaT, BATADAL, and WADI. Many attack and defense methods in these articles share operational principles with those used in computer vision.
L-BFGS [36] for generating adversarial samples tackles an intractable optimization problem. So, it simplifies this by solving a simpler optimization problem that utilizes a hyperparameter, sometimes challenging to select. The Fast Gradient Sign Method [37] (FGSM) resembles L-BFGS, albeit with the key distinction that FGSM generates a sample using a single iteration to create an adversarial sample. Projected Gradient Descent [38] (PGD) tackles the generation problem iteratively, introducing small perturbations at each step. Carlini–Wagner [39] (C&W) achieves a potent attack by solving a slightly different optimization problem, conceptually distinct from those derived from the L-BFGS method. Deepfool [40] is a hyperplane-based generation method, assuming a hyperplane exists between categories as a decision boundary. It calculates the distance from the current sample to each decision boundary, selecting the minimum distance as a perturbation added to the sample features.
@@ 2.3 Summary of the review
Since the main objective of the article is to create a benchmark, we pay particular attention to the most common methods described in the literature. This research focuses on analyzing vulnerabilities and implementing protection strategies within ASC. The following section elaborates on the mathematical aspects of the methods employed in this research.
@@ 3.2.1 Random noise
Random noise is a black-box attack based on adding random values to the input data:
@@ 3.2.2 Fast Gradient Sign Method (FGSM)
FGSM is a white-box attack based on the gradient of the loss function calculated for the input data. The signs of obtained gradient vector indicate the direction in which the input data should be changed to increase the probability of model error. The attack consists of shifting each value of the data by a step of size $\epsilon$, with a sign corresponding to the gradient:
@@ 3.2.3 FGSM distillation
Distillation can be used to create a black-box adversarial attack. Based on the input and output data of the model, a neural network classifier with an arbitrary architecture can be trained. Adversarial samples are obtained by attacking the resulting model by any white-box attack. In our study we used MLP architecture and FGSM attack.
@@ 3.2.4 Projected Gradient Descent (PGD)
PGD is an iterative modification of the FGSM white-box attack method. The main difference is that the data shift is done in several steps. After each step, the gradient signs are recalculated:
@@ 3.2.5 DeepFool
DeepFool is a white-box attack which minimizes the difference between the elements of the output vector $f(x)$ that correspond to the correct and incorrect fault type. Among all possible incorrect types, the closest in absolute value of the difference is selected. Minimization occurs in several steps, each defined as follows:
@@ 3.2.6 Carlini and Wagner (C&W)
C&W is a white-box attack that minimizes the sum of the shift value over the distance metric $D$ and the value of some auxiliary function $g$. Function $g$ takes negative values in case of incorrect classification. This optimization problem can be represented as:
@@ 3.3.1 Adversarial training
Adversarial training method consists of adding adversarial samples to the training set. The training loss function is given as follows:
@@
3.3.2 Autoencoder
Autoencoder can be used to reconstruct attacked data. During its training, the following loss function is minimized:
@@ 3.3.2 Autoencoder
3.3.3 Quantization
Quantization is a method that converts continuous values into some set of discrete values. This approach reduces the quality of the input data but can neutralize the impact of adversarial attacks. The fault diagnosis model must be retrained on quantized data.
3.3.4 Regularization
The fault diagnosis model can be protected by training using gradient regularization of the loss function over the input data:
@@ 3.3.4 Regularization
3.3.5 Distillation
Distillation defense method refers to the process of creating a copy of the original neural network model that is more resistant to adversarial attacks. The original neural network is called the teacher, and the new neural network is called the student. When teaching a student, so-called smooth labels are used, which are obtained using the activation function $\operatorname{softmax}(x,T)$ on the last layer of the teacher:
@@ 3.3.6 Combination
In our research we propose to look towards the joint use of several defense methods. Let’s imagine that there are methods $D_{1}$ and $D_{2}$ which, when protecting the model, do not greatly reduce its effectiveness. $D_{1}$ protects well against adversarial attacks of type $A_{1}$, and $D_{2}$ against attacks of type $A_{2}$. Then we can make an assumption that the $\text{combination}(D_{1},D_{2})$ will protect the model from both types of attacks $A_{1}$ and $A_{2}$. However, our experiments have shown that the behavior of this strategy is not always predictable. It takes many iterations to find the optimal settings for such combination. In our study, we explored combining adversarial training with a quantization defense method. During the training a data attacked first and then quantized.
@@ 3.4 Dataset
The Tennessee Eastman Process is a very popular dataset for benchmarking fault detection and diagnosis methods. It describes the operation of the chemical production line where the process smoothly flows from a normal state to a faulty. In our study we used extended by Reinartz et al. [43] TEP version that contains much more sensor data than the original one (5.2 Gb vs 58 Mb). It has 100 simulation runs for each of 28 fault types. Each run consists of 52 sensor values for 2000 timestamps. In our study, the input samples on which attacks will be carried out have the form of matrices $X^{k\times 52}$, where $k$ is the sliding window size.
@@ 4.2 Attacks on FDD models
At the next stage, unprotected models were attacked by six types of attacks with different $\epsilon$. For $\epsilon$ values, 20 points were selected in the range from 0 to 0.3 with a step of 0.015. Fig. 1 shows how the model’s accuracy degrades depending on the type and strength of the attack. It decreases significantly with small shifts in the attacked data for epsilon values less than 0.05.
@@ 4.3.1 Adversarial training
Figure 2: Accuracy of the TCN model protected by adversarial training with different settings: a) training on FGSM samples with fixed $\epsilon=0.1$; b) training on FGSM samples with set of $\epsilon$ values from the range (0, 0.3); c) training on PGD samples with fixed $\epsilon=0.1$; b) training on PGD samples with set of $\epsilon$ values from the range (0, 0.3).
@@ 4.3.5 Distillation
4.4 Combination of defense methods
The experimental results showed that various defense methods can be effective against some types of attack and not against the others. This fact suggests the idea of using several defense approaches together. In our study, we used a combination of adversarial training and quantization defense methods. When combining several types of protection, the final result is not always obvious and may turn out to be either better or worse than expected.
@@ 4.4 Combination of defense methods
Combining defense methods can be a very effective strategy for protecting against adversarial attacks. However, many iterations of experiments are required to select the optimal combination of methods and their parameters.
@@ 5 Discussion
Other methods such as quantization, regularization and distillation have shown high protection against some types of attacks and poor results against the others. These results prompted the idea of the possibility of combining them. In our study, we combined FGSM adversarial training and quantization defense method. This approach showed better results than these methods individually and is a promising direction for further research.
@@ Acknowledgment
Appendix A Additional tables and figures

### v3

@@ Abstract
Integrating machine learning into Automated Control Systems (ACS) enhances decision-making in industrial process management. One of the limitations to the widespread adoption of these technologies in industry is the vulnerability of neural networks to adversarial attacks. This study explores the threats in deploying deep learning models for Fault Detection and Diagnosis (FDD) in ACS using the Tennessee Eastman Process dataset. By evaluating three neural networks with different architectures, we subject them to six types of adversarial attacks and explore five different defense methods. Our results highlight the strong vulnerability of models to adversarial samples and the varying effectiveness of defense strategies. We also propose a new defense strategy based on combining adversarial training and data quantization. This research contributes several insights into securing machine learning within ACS, ensuring robust FDD in industrial processes.
@@ 1 Introduction
Automated Control Systems operate with a variety of digital and analog signals received from sensors and control mechanisms. An example is a chemical plant where a set of sensors reflect the condition of an industrial process. A common task is Fault Detection and Diagnosis (FDD), where one needs to predict and/or classify a failure based on sensors data. Such methods play a pivotal role in monitoring diverse industrial processes, ranging from chemical processes to electromechanical drive systems. The classification proposed by [1] categorizes these methods into three groups: those based on expert knowledge, mathematical models, and data-driven approaches. The latter one includes various approaches in machine learning, including neural networks [2, 3, 4]. Machine learning algorithms show themselves better than traditional methods based on rules and become more widespread in the area [5]. Recent studies demonstrate the success for FDD of various neural network architectures: Multi-Layer Perceptrons, Recurrent Neural Networks, Convolutional Neural Networks [6, 7].
However, another challenge appears: modern neural networks are vulnerable to adversarial attacks [8]. The idea is that the attacker slightly changes the input data (unnoticed) such that the FDD model prediction changes to incorrect. Such attacks are modeled in the literature [9, 10], but it remains unclear if there exist defense strategies good against a wide range of attacks. In order to address this challenge we benchmark attacks and defenses on the Tennessee Eastman Process dataset [11] where the task is the Fault Detection and Diagnosis (FDD) in a chemical process. We consider three various deep learning models — Multi-Layer Perceptron (MLP), model based on Gated Recurrent Units (GRU), and Temporal Convolutional Network (TCN). We subject these models to six different types of adversarial attacks and explore five defense methods. We analyze the success of these protective measures [^1] The source code to reproduce our results is available at https://github.com/AIRI-Institute/fdd-defense.. Then, a novel protection strategy is proposed, which employs several various defense methods.
The contributions of this work are as follows:
  • We benchmark popular attack and defense methods on the TEP dataset which shows that existing universal defense methods greatly reduce models quality on original data.
  • To address this issue we suggest a new defense approach based on adversarial training and data quantization and demonstrate its average effectiveness against various attacks.
  • We discuss benchmark results and conclude that autoencoders have a potential to be a universal defense methods but they need more research.
@@ 2.1 Classification of attacks on machine learning models
Attacks on machine learning models are typically categorized into several types, which vary based on the capabilities of the attacker in relation to the target and its characteristics throughout the model’s lifecycle. These attack types include evasion attacks [13, 14, 15], poison attacks [16, 17, 18] and exploratory attacks [19, 20, 21]. Additionally, some of these attack categories have further subgroups. Furthermore, attacks are segmented into three groups based on the level of information available about the model’s architecture and access to its internal parameters and/or requests: white-box, black-box, and gray-box attacks.
@@
Poison attacks [16, 17, 18] are a complex term in the literature. Typically, it refers to injecting poisoned samples into the training dataset with the aim of distorting the training process (so-called data poison attacks). Exploratory attacks involve sending queries to the model to understand its principles of operation. Such attacks can pursue various goals: stealing the model [22, 23], conducting membership inference attacks [19, 20, 21], and others. In this article the main focus was on evasion attacks, as they pose the greatest threat due to not requiring an insider attacker and directly impacting the model’s predictions.
In white-box attacks [24, 25], the attacker possesses complete information, enabling them to execute any operations on their instance of the deployed model (e.g., obtaining gradients, accessing output data from any layer) to construct perturbed samples. In the presence of defense mechanisms, these mechanisms are also susceptible to attack by the adversary.
Black-box attacks [26, 27, 28] assume that the attacker can only make a limited number of requests ($L$, where $1\leq L<\infty$) to the deployed model. The models provide the attacker with predictions such as label class or probability, semantic map segmentation, etc.
Gray box (or semi-white box) [29, 30] attacks represent an intermediate state between white-box and black-box scenarios. They involve the imposition of certain restrictions that provide some information about the learning model/process, albeit incomplete.
@@ 2.2 Most common methods of attack and defense
In the literature, there are numerous articles that explore issues akin to those addressed in this study. However, these methods are either examined on different datasets or pertain to disparate domain areas, necessitating adaptation for our purposes. Notably, many initial attacks were devised for image analysis models, and not all methods have been fully tailored for the domain area under our investigation. Therefore, in this review, we also consider these aspects. Articles such as [31, 32, 33, 34, 35] delve into the impact of attacks on various image datasets like CIFAR-10, CIFAR-100, ResNet-20, and MNIST. These articles also discuss various protective measures. Key attack methods include L-BFGS, FGSM, PGD, C&W, and DeepFool. While identifying the most prevalent defense methods can be challenging, several key approaches emerge, notably Defense-GAN and Adversarial Training. Moreover, articles such as [9, 36, 10, 37, 38, 39, 40] discuss attacks and defense methods on datasets like TEP and/or similar domain areas such as CARLA, Electra, SWaT, BATADAL, and WADI. Many attack and defense methods in these articles share operational principles with those used in computer vision.
In white-box adversarial attacks, access to gradients is a primary tool. Attackers exploit gradients by calculating the gradient of the loss function concerning the input. Then, they perturb the input in the direction of the gradient to maximize the loss. The mathematical details of the attack algorithms that will be used in this article are outlined in the next section (section 3).
The Fast Gradient Sign Method (FGSM), proposed by [41], generates adversarial examples with a single gradient step. It updates the input based on the direction of increasing loss, using a small multiple of the sign of the gradient. While FGSM is fast, its success rate for adversarial examples is low.
To improve the success rate of FGSM, [42] introduced the Projected Gradient Descent (PGD) method. Unlike FGSM, PGD takes multiple smaller steps in the gradient direction and clips the result by a specified value. Although PGD is more effective than FGSM in finding adversarial examples closer to the model’s decision boundary, it is computationally more expensive due to requiring multiple iterations.
For untargeted attacks, [43] proposed the DeepFool method optimized for the $L_{2}$ distance metric. It assumes the linearity of the decision boundary in neural networks and finds the minimum adversarial perturbation needed to fool the classifier. DeepFool iteratively identifies the direction that maximally changes the current prediction of the neural network and takes a small step in that direction until finding a true adversarial example.
A more sophisticated white-box attack, the C&W attack by [44], is applicable under various distance metrics: $L_{0}$, $L_{2}$, and $L_{\infty}$. This attack optimizes a loss function considering the distance between the original input and the adversarial example, along with the classifier’s prediction confidence. The optimization includes a constraint on the perturbation size, making the resulting adversarial example more realistic and challenging to detect.
@@
2.3 Interactions among defense methods
While the literature offers many different methods for defending machine learning models, there are few studies that explore building models combining multiple defense methods. In the paper [47] the authors examine the possibility of combining the most popular defense methods against evasion and poisoning attacks. The research concludes that many methods, at the level of algorithmic ideas, are incompatible, demonstrating this through practical examples. Therefore, constructing models that combine defense methods is a complex and underexplored task.
@@ 2.4 Summary of the review
Since the main objective of the article is to create a benchmark, we pay particular attention to the most common methods described in the literature. This research focuses on analyzing vulnerabilities and implementing protection strategies within ACS. The following section elaborates on the mathematical aspects of the methods employed in this research.
@@ 3.2.1 Random Noise
Random noise is the simplest black-box attack based on adding random values to the input data:
@@ 3.2.2 Fast Gradient Sign Method (FGSM)
FGSM [41] is a white-box attack based on the gradient of the loss function calculated for the input data. The signs of obtained gradient vector indicate the direction in which the input data should be changed to increase the probability of model error. The attack consists of shifting each value of the data by a step of size $\epsilon$, with a sign corresponding to the gradient:
@@ 3.2.3 FGSM Distillation
Distillation can be used to create a black-box adversarial attack as proposed in [58]. Based on the input and output data of the model, a neural network classifier with an arbitrary architecture can be trained. Adversarial samples are obtained by attacking the resulting model by any white-box attack. In our study we used MLP architecture and FGSM attack.
@@ 3.2.4 Projected Gradient Descent (PGD)
PGD [42] is an iterative modification of the FGSM white-box attack method. The main difference is that the data shift is done in several steps. After each step, the gradient signs are recalculated:
@@ 3.2.5 DeepFool
DeepFool [43] is a white-box attack which minimizes the difference between the elements of the output vector $f(x)$ that correspond to the correct and incorrect fault type. Among all possible incorrect types, the closest in absolute value of the difference is selected. Minimization occurs in several steps, each defined as follows:
@@ 3.2.6 Carlini and Wagner (C&W)
C&W [44] is a white-box attack that minimizes the sum of the shift value over the distance metric $D$ and the value of some auxiliary function $g$. Function $g$ takes negative values in case of incorrect classification. This optimization problem can be represented as:
@@ 3.3.1 Adversarial Training
Adversarial training method [41] consists of adding adversarial samples to the training set. The training loss function is given as follows:
@@
3.3.2 Defensive Autoencoder
Autoencoder can be used to reconstruct attacked data as proposed in [59]. During its training, the following loss function is minimized:
@@ 3.3.2 Defensive Autoencoder
3.3.3 Data Quantization
Quantization is a preprocessing method that converts continuous values into a set of discrete values on a uniform grid [60]. This approach reduces the quality of the input data but can neutralize the impact of adversarial attacks. The fault diagnosis model must be retrained on quantized data.
3.3.4 Gradient Regularization
The fault diagnosis model can be protected by training using gradient regularization [61] of the loss function over the input data:
@@ 3.3.4 Gradient Regularization
3.3.5 Defensive Distillation
Distillation defense method [62] refers to the process of creating a copy of the original neural network model that is more resistant to adversarial attacks. The original neural network is called the teacher, and the new neural network is called the student. When teaching a student, so-called smooth labels are used, which are obtained using the activation function $\operatorname{softmax}(x,T)$ on the last layer of the teacher:
@@ 3.3.6 Adversarial Training on Quantized Data
In this paper, we propose to use a combination of adversarial training and data quantization. As was shown in [60], quantization allows to clean the input from adversarial perturbation due to the grid alignment of discrete values. However, the size of the grid (quantization frequency), has an important role in this type of protection. If the grid is too wide, it reduces the quality of fault diagnosis, if the grid is too narrow, only a fraction of the data can be effectively recovered. On the other hand, adversarial training provides high model robustness but reduces the quality of diagnosis. This happens because during training, the data contains many adversarial examples that degrade the model’s ability to generalize important dependencies in the data that help diagnose faults. Thus, at high values of $\epsilon$ in adversarial training, the quality of the model drops significantly, otherwise it does not provide a sufficient level of protection.
We propose to use adversarial training on the data after quantization. Thus, during training, we attack the data with an adversarial attack such as FGSM. We then quantize this data and feed it into the input of the model as a training set. Quantization allows to reduce the strength of the attack, which in turn allows the model to generalize better during adversarial training. As a result, quantization helps the model to achieve better quality in adversarial training.
An additional advantage of this approach is that it does not require a separate model, as is the case with the distillation method or the autoencoder. It is also quite efficient in terms of computational time and memory, since quantization takes place in linear time and requires no additional memory, while adversarial training has the same complexity as training a model on the original data and also requires no additional memory.
@@ 3.4 Dataset
The Tennessee Eastman Process is a very popular dataset for benchmarking fault detection and diagnosis methods. It describes the operation of a chemical production line, where the process smoothly transitions from a normal state to a faulty one. In our study, we used a version of the TEP extended by Reinartz et al. [48] that contains significantly more sensor data than the original (5.2 Gb vs 58 Mb). This version includes 100 simulation runs for each of the 28 fault types. Each run consists of 52 sensor values for 2000 timestamps, and thus the input samples are in the form of matrices $X^{k\times 52}$, where $k$ is the sliding window size. All data in our experiments were standardized by removing the mean and scaling to unit variance.
@@ 4.2 Attacks on FDD models
At the next stage, unprotected models were attacked by six types of attacks with different $\epsilon$. For $\epsilon$ values, 20 points were selected in the range from 0 to 0.3 with a step of 0.015. We consider this range to be reasonable given that the data is scaled to a unit variance and the attack should not be detected by both human and automated defense mechanisms. Fig. 1 shows how the model’s accuracy degrades depending on the type and strength of the attack. It decreases significantly with small shifts in the attacked data for $\epsilon$ values less than 0.05.
@@ 4.3.1 Adversarial Training
Figure 2: Accuracy of the TCN model protected by adversarial training with different settings: a) training on FGSM samples with fixed $\epsilon=0.1$; b) training on FGSM samples with set of $\epsilon$ values from the range (0, 0.3); c) training on PGD samples with fixed $\epsilon=0.1$; d) training on PGD samples with set of $\epsilon$ values from the range (0, 0.3).
@@ 4.3.5 Defensive Distillation
4.3.6 Adversarial Training on Quantized Data
The experimental results showed that various defense methods can be effective against some types of attack and not against the others. This fact suggests the idea of using several defense approaches together. In our study, we used a combination of adversarial training and quantization defense methods.
@@ 5 Discussion
Other methods such as quantization, regularization and distillation have shown high protection against some types of attacks and poor results against the others. To address the limitations of individual defense methods, we explored the possibility of combining them. In our study, we combined FGSM adversarial training and quantization defense method. This approach provides good protection against most types of attacks (except for PGD adversarial examples with large $\epsilon$ values) with small losses in quality. It is computationally efficient and does not require additional memory. Other combinations of various defense methods can be explored in further research.
@@ Acknowledgment
APPENDIX A. Comparison of all methods